OpenStudy (anonymous):

A baseball coach is determining the batting order for the team. The team has 9 players, but the coach does not want the pitcher to be one of the first four to bat. How many batting orders are possible?

OpenStudy (dumbcow):

use the counting principle. For first 4 spots you have 8 players to choose from For last 5 spots you have 5 players to choose from \[\rightarrow \frac{8!}{(8-4)!}*\frac{5!}{(5-5)!} = \frac{8! *5!}{4!}\]

OpenStudy (anonymous):

so what would be the answer?

OpenStudy (dumbcow):

in terms of permutations its same as: \[8P4 * 5P5\]

OpenStudy (dumbcow):

the answer as far as a number you can figure out with a calculator :)

OpenStudy (anonymous):

how do I write it into the calculator

OpenStudy (dumbcow):

i dunno what kind of calc you are using....all scientific ones will have a factorial button (!)

OpenStudy (dumbcow):

or do it long way 8*7*6*5*5*4*3*2*1

OpenStudy (dumbcow):

i would think teacher would have gone over this before assigning this type of problem...hmm

OpenStudy (anonymous):

its not for me I haven't done this n 2 years

OpenStudy (dumbcow):

uh huh also use a computer calculator http://www.wolframalpha.com/input/?i=8%21*5%21%2F4%21

OpenStudy (anonymous):

this is for ma boyfriend cuz I had this 2 years ago n I honestly don't remember